how many pounds of peanuts costing $1.75 per pound must be combined with
cashews costing $3.50 per pound in order to obtain a 42-pound mixture that costs $2.50 per pound? Your answers may be decimals, but they must be exact.
Pounds of Peanuts = ___________________
Pounds of Cashews = ___________________

Answer by mananth(16946)

price ---------------- quantity
peanuts 1.75 ---------------- x pounds
cashew 3.50 ------ 42 - x pounds
Mixture 2.50 ---------------- 42
Total 42 pounds
1.75 x + 3.5 ( 42 - x ) = 105
1.75 x + 147 - 3.5 x = 105
1.75 x - 3.5 x = 105 - -147
-1.75 x = -42
/ -1.75
x = 24 pounds $1.75 peanuts
18 pounds $3.50 cashew

m.ananth@hotmail.ca